OpenStudy (anonymous):

Find to the nearest degree, all values of x in the interval 0°≤x<360° that satisfy the equation 3+tan²x=5tanx.

OpenStudy (campbell_st):

well rewrite the equation as \[\tan^2(x) -5\tan(x) + 3 = 0\] you have a quadratic equation, and looking at the coefficients you'll need the general quadratic formula to solve for tan(x) does that make sense

OpenStudy (anonymous):

Yeah that makes sense.. After that, I got 5+/-√12, and simplified the radical, then put them into two separate equations, 5+2√3/2 and 5-2√3/2... Now, what do I do?

OpenStudy (campbell_st):

ok... you now need to find the inverse tan values looking at \[\frac{5 + 2\sqrt{3}}{2}\] this is a positive angle so it exists in the 1st and 3rd quadrants.... so find \[x = \tan^{-1}( \frac{5 + 2 \sqrt{3}}{2})\] 1st quadrant will be the answer for x 3rd quadrant will be 180 + x you need to repeat the process for the other value

OpenStudy (campbell_st):

you will end up with 4 angles.... 2 in the 1st quadrant and 2 in the 2nd quadrant as you need to find \[x = \tan^{-1}(4.23)...... and..... x = \tan^{-1}(0.7679)\]

OpenStudy (anonymous):

But first, do I put the whole tan inverse in the calculator like this: \[\tan^{-1}\](5+2√3/2) or \[\tan^{-1}\](5+2√3)/2 ? Does that make sense?

OpenStudy (campbell_st):

oops should be 2 angles in the 3rd quadrant

OpenStudy (campbell_st):

yep thats it.... just as easy to make it a decimal value.... since you are using inverse tan.... to find angles

OpenStudy (anonymous):

Thank you so much, I found the answers to the problem, you're a major help!! Again, thank you!

OpenStudy (campbell_st):

well done